Calculate VAT

goods are sold at R3690 calculate VAT


1
Expert's answer
2021-09-17T17:42:23-0400

Solution:

The standard VAT rate for South Africa is 15"\%" for goods and services.

The amount of R3690 is inclusive of 15"\%" VAT.

Derive the gross amount:

115"\%" = 3690

100"\%" =?


= "\frac{100\% \times 3690}{115} = 3,208.70"


The gross amount of the goods exclusive of VAT = 3,208.70


VAT = Vat rate "\times" Gross Amount = 15"\%\times"3,208.70 = 481.30


VAT of the goods  = 481.30